OpenStudy (anonymous):

Hey can you guys explain to me how to find the exact volume of a cylinder? You take the radius and split it in half and times it by its self, times that by the height and thats it? Or do you divide by 3.14 or times it by 3.14? Just that last step with pi I'm not sure what to do. Thanks! :)

OpenStudy (anonymous):

You take the diameter and split that in half to find the radius. To find the volume of a prism you need to find the area of the 'face' and multiply that by the depth/height. i.e. the face is the circle pi*r^2 and multiply that by the height

OpenStudy (anonymous):

You multiply 3.14 by the radius and the height

OpenStudy (anonymous):

So say the height of a cylinder was 8 and the diameter was 4. It would be, 1. Half of 4, 2. 2. 2x2= 4 3. 4 x 3.14= 12.56 4. 12.56 x 8 = 100.48 5 and thats it?

OpenStudy (phi):

if you want *exact* use the symbol \( \pi\) rather than 3.14 (pi is only approximately 3.14) so volume= \( \pi r^2 h = \pi\ 2^2 \ 8 = 32 \pi \)

OpenStudy (phi):

you might recognize \(\pi r^2\) is the formula for the area of a circle (the area of the bottom of the cylinder)

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

Its really easy once you get how to do it